OpenStudy (anonymous):

IN THE PRODUCTS 2a*1a = 41a, 2b*1B=37b, a AND b REPRESENT POSITIVE INTEGERS. THEN a+b

OpenStudy (anonymous):

is this a question?

OpenStudy (anonymous):

if it is ill help you

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

2a * la = 4*la 2b*1b=37b ????

OpenStudy (anonymous):

?????? >-<

OpenStudy (anonymous):

thought you just replied haha lol

OpenStudy (anonymous):

2b*1b=37b

OpenStudy (anonymous):

yes